greg is on a 2400 kcal/day diet. he ate a breakfast that contained 80 g of fat for breakfast. approximately how much of his daily amdr for lipid has he consumed?
Answer #1

1 gram of lipid (fat) provides 9 Kcal

Therefore, number of Kcal provided by 80g fat = 80 * 9 = 720 Kcal

So, AMDR for lipid consumed = (720 / 2400 ) * 100 = 30%.